# Representational Alignment Supports Effective Machine Teaching

### v1

@@ Abstract
A good teacher should not only be knowledgeable; but should be able to communicate in a way that the student understands – to share the student’s representation of the world. In this work, we integrate insights from machine teaching and pragmatic communication with the burgeoning literature on representational alignment to characterize a utility curve defining a relationship between representational alignment and teacher capability for promoting student learning. To explore the characteristics of this utility curve, we design a supervised learning environment that disentangles representational alignment from teacher accuracy. We conduct extensive computational experiments with machines teaching machines, complemented by a series of experiments in which machines teach humans. Drawing on our findings that improved representational alignment with a student improves student learning outcomes (i.e., task accuracy), we design a classroom matching procedure that assigns students to teachers based on the utility curve. If we are to design effective machine teachers, it is not enough to build teachers that are accurate – we want teachers that can align, representationally, to their students too.
@@ 1 Introduction
What makes an effective teacher? Massive Open Online Courses (MOOCs) promised to revolutionize education by having educators from top research institutions record lectures in their domain of expertise and make course materials available online for any learner to interact with. However, after many trials, this expert-first approach to online learning proved not as generally effective and accessible as hoped for [1], with courses delivered more traditionally by local teachers often showing better outcomes, in-person and online [2].
Similarly, many can relate to having teachers who are brilliant experts in their domains, but who struggle to instill knowledge in their students [3, 4]. These two pedagogical phenomena share a common thread – for their teaching to be effective, teachers should not only be knowledgeable in their domain but should also be able to appropriately map what they know into a space accessible to their audience of learners. As powerful AI systems become increasingly prevalent in society, it becomes increasingly pressing for us to understand how these systems can be turned into effective teachers for their human users – thought partners that enhance us rather than supplant us. However, as these systems begin to outperform humans on some tasks [5, 6, 7], their internal representations, at times, seem to become less human-like [8, 9], setting up a tension between domain expertise and the ability to map knowledge into human-understandable spaces.
In this work, we bridge ideas from the burgeoning subfield of representational alignment [10] with work on machine teaching, and the cognitive science of pedagogy, beginning to address the question: is a teacher who is an expert but representationally misaligned from a student better than a teacher who is less of an expert but more representationally aligned? We hypothesize that there is a tradeoff between expertise and representational alignment. Clarifying this tradeoff is important in practice, as there may be design pressures in building machine teachers wherein practitioners need to choose between increasing accuracy versus increasing alignment. Through simulations, we trace out a utility curve expressing this tradeoff. We validate our curve through experiments in which machines teach humans, demonstrating that expert (machine) teachers may not be as valuable as peer (machine) teachers who are representationally aligned to their students.
Figure 1: Schematic of various teaching set-ups. (Left:) Dyadic interaction between a single teacher and a single student. The teacher selects an example, or set of examples, to communicate to the student. Teachers be “student-centric” and represent their student when deciding which examples to select (left bottom). Middle: “‘Classroom” setting wherein a single teacher selects a set of examples to provide to all students in their “class” at once; student-centric teachers reason about all students in their classroom to make their selections (middle bottom). Right: “School” setting wherein many different teachers are matched with different students; each student is matched with a single teacher.
Alignment is not necessarily a fixed quantity of a teacher. Skilled human teachers can often infer their students’ representations and adapt their teaching accordingly. To compare such “student-centric” teachers to the “self-centered” ones discussed above, we extended our analysis to “classroom” settings where a single teacher has to teach multiple students using the same set of examples (see Figure 1). While we find that, as expected, student-centered teachers have better teaching outcomes than self-centered teachers, we show that they are nonetheless impacted by the representational alignment of their students with each other, and that the average student performance drops as class size increases. This means that simply assigning all students in a school to the most expert student-centered teacher is not always an optimal strategy for maximizing student group outcomes.
Combining these insights, we leverage our utility curve to design a classroom matching procedure that, when given a pool of students and teachers (i.e., a “school”), strives to maximize student outcomes by assigning students to teachers based on their accuracy and representational alignment. We test our proposed algorithm with various learning outcome metrics used in real-world settings including average performance, worst-case performance , and passing rate. We show that our matching algorithm significantly improves learning outcomes for students compared to both random matching and standardized curricular materials (e.g., a recorded MOOC lecture distributed to all students). Furthermore, we find that having student-centered teachers available at a school, even if these teachers have very little expertise in the domain being taught, can greatly improve student outcomes, especially for students who are representationally unique relative to the other students and teachers. Our proposed framework and experimental results provide early insights that could inform the general study of pedagogy and human learning, as well as the specific problem of designing effective machine teachers in an age of models trained from human feedback (i.e., reinforcement learning with human feedback [11]).
2 Problem Setting
@@
Figure 2 shows a schematic of our teaching and representation alignment framework. Consider a space $X$ of stimuli. We consider the case where the teacher tries to teach the students some function $f:X\rightarrow C$. We illustrate a simple case in Figure 2 wherein $C$ is a binary classification $C=\{0,1\}$ dividing $X$ into two regions ($C=0$ and $C=1$ are represented in Fig. 2 in light and dark gray, respectively). The teacher observes label function $f^{\prime}:X\rightarrow C$, which may be different from $f$.
@@
3 Simulations
3.1 Setup
Task. We construct a grid environment of size $N\times N$ where there are $K$ categories. Each grid cell is defined by its location within the grid with a label $k\in\{1...,K\}$. The student’s goal is to correctly label each point. To support the student, the teacher can reveal labels on the grid (i.e., creating teaching materials $L_{0}$ which consist of grid points and their associated labels). We consider two different label structures ($f$): one wherein there are $N$ classes (one class per column) and one with 4 classes (partitioning the grid into quadrants). Our experiments are reproducible and all the implementations for all computational experiments will be made available open-source upon publication.
Representation. The coordinates of grid cells form the student ($Y_{s}$) and teacher’s ($X$) representations. We compute representational alignment as correlation between pairwise Euclidean distances. We corrupt the teacher’s representation (in order to create misalignment between teacher space $X$ and student space $Y_{s}$) by sampling a number of grid cells to swap such that cell $(i,j)$ for the student appears as cell $(i^{\prime},j^{\prime})$ for the teacher (i.e., the mapping $T_{s}$ from the teacher to the student would reverse these swaps).

### v2

# Representational Alignment Supports Effective Teaching

@@ Abstract
A good teacher should not only be knowledgeable, but should also be able to communicate in a way that the student understands – to share the student’s representation of the world. In this work, we introduce a new controlled experimental setting, GRADE, to study pedagogy and representational alignment. We use GRADE through a series of machine-machine and machine-human teaching experiments to characterize a utility curve defining a relationship between representational alignment, teacher expertise, and student learning outcomes. We find that improved representational alignment with a student improves student learning outcomes (i.e., task accuracy), but that this effect is moderated by the size and representational diversity of the class being taught. We use these insights to design a preliminary classroom matching procedure, GRADE-Match, that optimizes the assignment of students to teachers. When designing machine teachers, our results suggest that it is important to focus not only on accuracy, but also on representational alignment with human learners.
@@ 1 Introduction
The proliferation of digital education resources and AI systems has enabled human and machine teachers to reach potentially millions of students. For example, Massive Open Online Courses (MOOCs) promised to revolutionize education by having top educators record lectures in their domain of expertise and make course materials widely accessible online for many learners. However, this expert-first approach to online learning was not as effective and accessible as hoped for (Reich and Ruipérez-Valiente, 2019), with courses delivered by local teachers often showing better outcomes, in-person and online (Kelly, 2014). More recently, AI systems like ChatGPT have gained hundreds of millions of users, many of whom are using them, or the educational applications they power, to learn new subjects. While these systems can now outperform humans on some tasks (Strachan et al., 2023; Van Veen et al., ; Thirunavukarasu et al., 2024), their internal representations are not often human-like (Fel et al., 2022; Muttenthaler et al., 2022), highlighting the distinction between domain expertise and the ability to map knowledge into human-understandable spaces. This tension is neither new nor unique to AI; professors can also be experts in their fields that struggle to communicate knowledge to students (Carter et al., 1987; Hinds et al., 2001). Yet many recent public education proposals explicitly focus on increasing teachers’ domain expertise (e.g., Ontario, 2024), and much AI research continues to focus on improving the expertise of the agents being developed. Understanding further factors of effective teaching can help determine strategies for improving outcomes in classrooms.
We aim to bring together ideas from the burgeoning subfield of representational alignment (Sucholutsky et al., 2023b), machine teaching, and the cognitive science of pedagogy to shed light on further improvements for classrooms. We propose that 1) representational alignment between teachers and students, and 2) the size and diversity of the classroom, are both critical for determining the effectiveness of human and machine teaching (Figure 1B). To test this hypothesis, we design a simple modular student-teacher cognitive task environment called “Grid Manipulation of Representational Alignment and Domain Expertise” (GRADE) that enables the experimenter to independently control the teacher’s expertise on the task, and the degree to which their representations of the task are similar to the student’s (Figure 1A). Through simulations and a study where machines teach humans, we establish the relationship between teacher expertise, teacher-student alignment, and student performance. We find that representationally aligned teachers with a high error rate on the underlying task can outperform highly accurate but representationally misaligned teachers (Figure 1F). These results suggest that if a teacher adapts their representations to match the student, then the student’s learning outcomes can significantly improve. We then extend our task from a teacher interacting with individual students to interacting with a class of representationally diverse students where the material they present is broadcasted to all students in the class (e.g., a lecture; see Figure 1D) and determine that class size and representational diversity moderate the effect of representational alignment on student outcomes (Figure 1G). Finally, we design a preliminary classroom matching procedure, GRADE-Match, that takes into account representational alignment, teacher expertise, and class size to optimize learning outcomes when assigning students to teachers (Figure 1E). We find that it outperforms both random assignment and MOOC-style assignment (Figure 1H). Our study emphasizes the importance of considering student-teacher representation alignment – not just teacher expertise – in pedagogical settings. This is especially important for designing AI thought partners that can think with us to help us grow (Collins et al., 2024) and tools that help personalize suggestions to individual students (Wang et al., 2024).
2 GRADE: Grid Manipulation of Representational Alignment and Domain Expertise
We designed a new controlled task domain, called GRADE, where stimuli are arranged on a grid and labeled. Teachers know all the labels for the stimuli, though their expertise can be modulated by corrupting the true labels to get teachers with varying accuracy. Students do not yet know the labels but do see the stimuli. We focus on the setting where the teacher selects some labeled examples to reveal to the student. The arrangement of the stimuli on the grid might vary between the student and teacher, allowing us to manipulate representational alignment. We show an example of misaligned grids with two labeled examples chosen by the teacher in Figure 1A. GRADE lets us use any stimuli that can be arranged on a grid (e.g., based on pre-set features, as in the “salient-dinos” case in our human experiments, or even amortized embeddings). Here, we define representation alignment with respect to stimuli locations on the student and teacher grids; that is, we compute the Euclidean distance between pairwise swaps between stimuli. GRADE permits modularly-specifiable representation functions; we refer to  Sucholutsky and Griffiths for a survey of a myriad of ways of measuring representation alignment. Additionally, we focus on a one-shot case where the teacher makes one round of selections. However, researchers can easily extend GRADE to multi-turn interactions. Appendix C contains a theoretical formalism of our setting.
Figure 1: Overview. A: GRADE task; teacher and student receive misaligned grids (numbers only represent re-arranged elements, participants do not see them). Teacher is shown all labels (shown as colors) and reveals one per class to the student. Teacher’s error rate is controlled by mislabeling their grid. B: Our hypothesized causal model. C: Dyadic interaction between a teacher and a student. “Student-centric” teachers infer the student’s representations, making them fully aligned. D: “Classroom” setting where teacher broadcasts examples to all students (who have individual differences in representations); student-centric teachers jointly optimize over all students in the class. E: “School” setting where teachers are matched with students; each student is matched with a single teacher. F: Utility curve relating teacher error rate, representational alignment, and student accuracy in simulations. G: Average accuracy and standard errors in a student-centric class as a function of class size in simulations. H: Average accuracy and standard errors across a school achieved by different matching procedures in simulations.
3 RQ1: Does representational alignment affect teaching outcomes?
We begin to explore the relationship between representation alignment and student outcomes in two settings. We instantiate GRADE with two kinds of stimuli that can be arranged on an $N\times N$ grid with $K$ underlying classes: simple-features (where each stimulus is only represented by its $(x,y)$ coordinates; see Figure 7) and salient-dinos (stick figure images with features varying based on grid location). First, we simulate student-teacher interactions with simple 1-NN agents (see Appendix  E.1) in the simple-features setting. We then generalize these findings with real human learners in both tasks. We include details of our teacher and student models, as well as our human experiment, in Appendix E.3.
3.1 Representational (mis)alignment in simulations
In Figure 1F, we trace out a relationship between student-teacher representational alignment, teacher error rate, and student accuracy. We uncover instances wherein students can achieve higher performance by learning from teachers who are more erroneous (“less expert”) provided the teachers are representationally aligned with the students than comparatively more expert but representationally misaligned teachers, underscoring that it is not just the accuracy of a teacher that matters for student learning outcomes. For a fixed teacher error rate, higher representational alignment is always better for a student (provided the error rate is not too high). We uncover similar curves across grid sizes and the number of categories (see Appendix E.2).
3.2 Representational alignment of machine teachers and human students
We then generalize our findings through human experiments with $N=480$ participants (see Appendix  E.3). We construct a utility curve paralleling our simulations by post-hoc varying teacher error rate (see Appendix G.3). We find in Figure 2 that across both tasks (simple-features and salient-dinos), generally, higher representational alignment induces higher average student accuracy, and report correlations in Table 1. We find that even large increases in teacher error rate can be offset by increasing representational alignment (e.g., a teacher with error rate 00 and representational alignment of $0.3$, has similar student outcomes as a teacher with error rate $0.4$ and representational alignment $0.8$). However, we note that the ordering of high representational alignment is less clear, particularly for the settings where each class corresponds to a column. We posit that people have a strong prior against classes being distributed as columns, and find that especially for the column conditions, participants would often label using strategies that did not correspond to nearest neighbor classification (e.g., several participants labeled in a way that corresponded to different types of tilings).
Figure 2: Relating teacher error rate and representational alignment between machine teachers and human students to student accuracy. From left to right: simple-features one class per quadrant; salient-dinos one per quadrant; simple-features one class per column (6); salient-dinos one class per column (7). Format follows Figure 1F.
4 RQ2: How does class size and student representational diversity moderate the effect of representational alignment on student outcomes?
We have demonstrated that both a teacher’s representational alignment and their accuracy matter for student outcomes. So far, our teachers have been self-centered; they use a single representation to select labeled examples. This approach suits machine teachers unable to adapt to specific students but does not capture capabilities of adaptable human or future machine teachers. Additionally, classroom teachers often address multiple students with differing representations, making example selection more complex. Here, we consider student-centric teachers who aim to maximize the average performance of a student pool by simulating likely student learning outcomes to various selections in an “inner loop” optimization (see Appendix F.6). While our earlier findings suggest student-centric teachers may enhance learning by becoming representationally aligned with students, we hypothesize that this effect will be moderated by the group’s size and representational diversity, since teachers must optimize for all students simultaneously.
4.1 Setup
We extend GRADE to investigate classrooms of varying sizes. Because we sample students for each class from the same pool of representationally diverse students (Appendix  F.1), increasing classroom size will generally increase representational diversity.
4.2 Results
We investigate the relationship between classroom size and student performance by sampling teachers with a range of error rates (between 00 and $0.5$ in increments of $0.1$) and classroom sizes ($10$ seeds per setting). Each student-centered teacher optimizes for their class through $T=100$ inner loop iterations. We then marginalize over our sampled error rates to compute an expected average classroom accuracy per classroom size. We find that the performance of students in a classroom with a student-centric teacher is initially high (i.e., with a class of only a single student, the student-centric teacher would be equivalent to a fully representationally aligned teacher in the dyad setting) but falls off rapidly as a function of classroom size and then plateaus (as shown in Figure 1G).
5 RQ3: Can we match teachers and students to improve outcomes?
Given a “school” of teachers and students, how can we simultaneously group students into “classrooms” and determine which teacher to allot to which class? We begin to explore this question through a series of “classroom matching” experiments. We develop a classroom matching procedure, GRADE-Match, which given a pool of students and teachers, assigns groups of students to teachers based on our representational alignment-teacher utility curve. We emphasize though, that our analogy to “classrooms” and “schools” is explored in simulation with machines teaching machines; substantial future work is required to investigate the generalization of possible links between representational alignment, teacher error rate, and classroom properties in practice.
5.1 Setup
Student and teacher pools. We focus on our simple-features setting and extend our dyad (single teacher, single student) setting to simulated pools of teachers and students over our same $6\times 6$ grid. We design two different pools of students and teachers (unstructured and structured). We include pool construction and generalization experiments to the salient-dinos setting in Appendix  F.1 and  F.5.
Matching procedures. We propose matching students using our utility curve to estimate their accuracy (Grade-Match (Ours)). We compute the representational alignment between a student and teacher and index into a bucketed version of the utility curve [^1] We recompute the curve by also averaging over samples of corrupted students as our first utility curve was constructed for dyad setting wherein the student’s grid was never corrupted (see Appendix F.2). that we construct in Section 2 using both the representational alignment and teacher’s expected error rate (which we assume we have access to). The resulting metric is the student’s expected performance under a specific teacher and classroom. We iterate over all teachers for each student and select the teacher who helps the student achieve the highest expected performance. We consider three baselines: (i) Random matching of students to teachers, (ii) MOOC which matches all students to the lowest error rate teacher, and (iii) Optimal wherein we use a brute force-search to match students to the highest attainable accuracy, giving an indication of the upper limit of performance that a matching algorithm could possibly achieve. Gaps between (ii) and (iii) further drive home the importance of going beyond teacher accuracy when pairing students to teachers.
5.2 Results
Our matching algorithm, which groups students to teachers based on their representational alignment and teacher error rate, generally outperforms random matching and, particularly for top-performing students, is better than having assigned the student to an expert (minimal error rate teacher; MOOC) who may be representationally distinct (see Figure 1H and Appendix Tables 2 and 3). This observation is intriguing – students may not achieve their full potential when paired with a representationally misaligned teacher, even if that teacher is an expert. We observe performance gains for our utility curve-based matching across both pool types. However, we do not yet attain optimal matching performance, perhaps due to a mismatch in our utility curve.
6 Discussion and limitations
Expertise on a task is not sufficient to be a good teacher; representational alignment matters too. Using a new controlled experimental paradigm (GRADE), we trace out a utility curve between teacher accuracy, teacher-student representational alignment, and student accuracy to characterize the crucial relationship between representational (mis)alignment and student learning outcomes. We put this utility curve to work to better match teachers to students based on representational alignment. Our work underscores the importance of teachers representing a diversity of students and arranging student-teacher groups to ensure there is at least one teacher that any student can effectively learn from. This motivates further investigation into representational alignment and its influence on pedagogical effectiveness in multiple learning settings, like teacher-student interactions and peer mentorship.
Yet, we emphasize that our work is a first step in the study of the relationship between representational alignment, teacher efficacy, and student-teacher matching. Our simulations always assume that students are 1-NN classifiers, which grossly undercuts the richness of human behavior. Further, our simulated students’ representations are fixed; in practice, students adapt their representations over time. We also only consider single-turn, single-lesson settings, wherein students have no indication of the reliability of the teacher. We look forward to investigations that leverage and extend GRADE to go beyond our simple yet revealing initial setting.
7 Acknowledgments
We thank Alexis Ross, Dilip Arumugam, and Kartik Chandra for helpful conversations that informed this work. KMC gratefully acknowledges support from the Marshall Commission and the Cambridge Trust. AW acknowledges support from a Turing AI Fellowship under grant EP/V025279/1 and the Leverhulme Trust via CFI. This work was supported by a grant from NOMIS Foundation to TLG. MKH was supported by NSF Award #2434192
Appendix A Broader impact and societal risks
As we discuss in Section 6, our work portends broader implications for the design of machine-human teaching setups where machines are intentionally built with representation alignment in mind, as well as representation diversity to safeguard against threats to inclusivity. It is possible that our simulations could support interventions in real classrooms, e.g., informing classroom size decisions drawing on measures of the representational diversity of a classroom pool and the expertise of the teacher. However, we heed caution in over-generalizing our results to settings where real student experiences and learning potential is at stake. Broad-brush application of AI systems in education has not been met with universal success (Reich, 2020) and inappropriately incorporated can have unintended impacts on student success (Fischer et al., 2020).
Appendix B Related work
Learning Sciences. The extended learning sciences community has studied aspects of what makes for a good teacher or computer-based teaching system. The expertise or quality of the teacher with respect to excellence of schooling, certification, and a teacher’s own test scores have been observed to positively affect student learning (Rice, 2003). More classroom-adaptive qualities, like a teacher’s amount of experience in classrooms and teaching strategies employed (i.e., pedagogy) are also top contributing attributes (Rice, 2003). Closeness of representation to students with respect to demographic features has been shown to lead to more effective student performance (Dee, 2004), in part due to the role model effect, but also because teachers closer in these dimensions can serve as sociocultural interlocutors, helping translate the relevance of material to students (Egalite et al., 2015; Harfitt, 2018). Intelligent Tutoring Systems (Anderson et al., 1985), growing out of the cognitive and learning sciences, have been a consistently effective paradigm of computer-based teaching (Wang et al., 2023), primarily utilizing the pedagogy of mastery learning (Bloom, 1984). They adapt the amount of prescribed practice based on a representation of the student’s level of mastery of the skill being worked on and provide procedural remediation in the problem-solving context. In a two-year, large-scale evaluation, a commercial ITS was found to be effective overall, but only demonstrated superior learning gains to standard classroom instruction in the second year. It was hypothesized that this may have been due to teachers needing to learn how best to align their classroom to the technology (Pane et al., 2014).
Machine teaching. Machine teaching aims to study the problem of teaching efficiency by characterizing such efficiency as the minimal number of effective data examples that is needed for a learner to learn some target concept. It has an intrinsic connection to optimal education (Zhu, 2015), curriculum learning (Liu et al., 2017; Korakakis and Vlachos, 2023) and optimal control (Lessard et al., 2019). Depending on the type of learner, machine teaching can be performed in a batch setting (Zhu, 2015; Zhu et al., 2018; Liu and Zhu, 2016) or an iterative setting (Liu et al., 2017, 2018, 2021; Qiu et al., 2023; Zhang et al., 2023). The batch teaching aims to find a training dataset of minimal size such that a learner can learn a target concept based on this minimal dataset. The iterative teaching seeks a sequence of data such that the learner can sequentially learn the target concept within minimal iterations. Complementary to these works, our findings indicate that, alongside the quality of examples that the teacher selects, it is also critical for both the teacher and the student to share similar representations.
Pragmatic communication. Successful communication rests on our ability to understand others’ beliefs and intentions (Gweon, 2021; Vélez et al., 2023). Indeed, even young children are sensitive to others’ knowledge and competence when teaching (Liszkowski et al., 2008; Bridgers et al., 2020) and learning (Bass et al., 2022; Csibra and Gergely, 2009; Bonawitz et al., 2011) from others. Inspired by Gricean pragmatics (Grice, 1975), recent computational models have formalized this process as recursive reasoning about others’ latent mental states (Chen et al., 2022; Goodman and Frank, 2016; Shafto et al., 2014). Such pragmatic models have been used to study and facilitate human-AI interaction  (Sumers et al., 2021, 2022; Lin et al., 2022; Andreas and Klein, 2016; Dale and Reiter, 1995; Fried et al., 2018; Wang et al., 2016, 2020; Ho et al., 2016; Zhi-Xuan et al., 2024; Liu et al., 2024). Crucially, however, when either party fails to accurately model the other’s beliefs or perspective, human-human (Aboody et al., 2023; Sumers et al., 2023) and human-AI (Milli and Dragan, 2020; Sumers et al., 2022) communication can be significantly degraded. Our work adds to this literature by formalizing and analyzing the effect of representational misalignment on communication.
Representational alignment. Representational alignment (Sucholutsky et al., 2023b) offers a conceptual and grounded mathematical framework for characterizing teaching settings wherein two or more agents engage on some task. Already, ideas from representational alignment are providing new ways of thinking about machine learning efficiency  (Sucholutsky et al., 2023a; Sucholutsky and Griffiths, 2023), value alignment  (Rane et al., 2023; Wynn et al., 2023), disagreement (Oktar et al., 2023), and applications like human & machine translation and conversation (Niedermann et al., 2024; Huang et al., 2024). In this study, we show that representational alignment is a key dimension in predicting and optimizing student outcomes, with similar importance as the teacher’s subject expertise.
Appendix C Theoretical formulation underlying GRADE
@@
We offer a deeper theoretical formalism for our setting. Figure 3 shows a schematic of our teaching and representation alignment framework. Consider a space $X$ of stimuli. We consider the case where the teacher tries to teach the students some function $f:X\rightarrow C$. We illustrate a simple case in Figure 3 wherein $C$ is a binary classification $C=\{0,1\}$ dividing $X$ into two regions ($C=0$ and $C=1$ are represented in Fig. 3 in light and dark gray, respectively). The teacher observes label function $f^{\prime}:X\rightarrow C$, which may be different from $f$.
@@
Appendix D Compute details
All experiments were run on an 8-core, 16 GB memory laptop. Experiments were run exclusively on CPUs and were all runnable within at most three hours. Our experiments are reproducible and all the implementations for all computational experiments will be made available open-source upon publication.
Appendix E Additional details on task setup for the single teacher-single student setting
E.1 Student and teacher models